# A theory of quasiballistic spin transport

### v2

@@ I Introduction
Lattice models in physics generically exhibit diffusive transport of their local conserved charges at high temperature. The most commonly encountered exceptions to this rule arise in one spatial dimension, as a consequence of integrability. At one extreme, there exist models that can be mapped to non-interacting particles, such as the spin-$1/2$ XX [1] and Haldane-Shastry [2] chains, which therefore exhibit ballistic transport of all their local conserved charges [3, 4, 5]. A more complicated intermediate behavior arises in the spin-$1/2$ XXZ chain, which can support ballistic, diffusive, or superdiffusive transport of spin, depending on the model anisotropy [6].
@@
Thus, unlike in continuous systems, faster-than-diffusive transport in lattice models is generally linked to the presence of integrability [7, 4]. As one perturbs away from a given integrable point, such faster-than-diffusive dynamics usually gives way to normal diffusion. However, the relaxation timescales controlling this crossover to normal diffusion can be extremely long, even when the system under consideration is chaotic in all other respects [8, 9, 10, 11, 12, 13, 14, 15, 16].
@@
In particular, several recent works have observed numerical evidence for long-lived and transiently ballistic (or “quasiballistic”) dynamics of conserved charges in systems that appear to be far from integrability. For example, spin transport in power-law-interacting spin-$1/2$ XXZ chains [14, 15] (see also [8]) and charge transport in interacting fermionic lattice models in any spatial dimension [16] can exhibit this phenomenon. Both sets of work focus on infinite temperature, where such behaviour is intuitively least likely to occur. In the fermionic case, this physics was explained theoretically by computing an effective lifetime for the charge current, which could be made arbitrarily small as a function of the interaction range. Meanwhile, no detailed theoretical explanation has been proposed for quasiballistic spin transport in general and providing such an explanation is the primary goal of this work.
@@
The specific findings to be explained are as follows. It has been argued [14, 15] that for spin-$1/2$ XXZ chains with power-law exchange interactions $J(r)\propto 1/r^{\alpha}$ between spins, long-lived and transiently ballistic modes of spin can arise for all values of the power-law exponent $\alpha$, with maximal enhancement of these ballistic modes for model anisotropies $\Delta^{*}(\alpha)\approx e^{2-\alpha}$ interpolating smoothly between the Haldane-Shastry chain (at $\alpha=2$) and the XX chain (as $\alpha\to\infty$), both of which exhibit ballistic spin transport as discussed above. However, the evidence that this locus exists and the extrapolation of its shape have so far been purely numerical. Our main contribution in this paper is an analytical prediction for the locus $\Delta^{*}(\alpha)$ on which the lifetime of the spin current is maximized, as a ratio of two infinite series. A nontrivial consistency check on our prediction is that it interpolates smoothly between the two ballistic points. However, our expression has the asymptotic behavior $\Delta^{*}(\alpha)\sim 3\cdot 2^{-\alpha}$ as $\alpha\to\infty$, whose exponential dependence on $\alpha$ matches the earlier conjecture from numerics [14, 15], but decays more slowly when $\alpha$ is large.
@@
We emphasize that Hamiltonians of this type are not merely a theoretical curiosity. Recent advances in quantum devices have renewed interest in the dynamics of long-range interacting spin chains [17, 18, 19, 20, 21]. Importantly, these platforms inherently favour the study of real-time dynamics. Moreover, their relatively high effective temperatures, which can be undesirable when it comes to exploring traditional questions about ground-state physics, are ideal for studying the far-from-equilibrium transport properties of interest in this work. There now exist several concrete proposals [22, 23, 24] for how to realize precisely the power-law interacting XXZ Hamiltonians studied here using trapped ions and a recent experiment [25] probing their dynamics. It should, however, be borne in mind that relatively small power-law exponents $0\leq\alpha\leq 3$ are most physically natural in these platforms [26], while the discussion in our paper applies to $\alpha>3/2$.
@@
In this work, we pursue a less ambitious yet analytically tractable version of the path sketched above. Rather than ruling out ballistic spin transport fully for $\alpha\neq 2$, we rule out the possibility of ballistic spin transport protected by a strictly two-body conservation law, as it is at the Haldane-Shastry [27, 28, 4] and XX points. Similarly, rather than attempting to evaluate the spin diffusion constant from the Kubo formula, we estimate the lifetime of the spin current directly from its instantaneous rate of change, a strategy that has proved effective in fermionic systems [16] (see also [29]) and is reasonable in the absence of ballistic transport. We conclude by discussing the scope of our approach and the connection between integrability and transport more broadly. In particular, we prove that the only spin-$1/2$ chains of the form Eq. (1) with an exactly conserved spin current are Haldane-Shastry and XX chains (Theorem 1 of Appendix A.1).
@@ II Decay rate of the spin current
where $a(r)$ and $b(r)$ are arbitrary even functions and it will be convenient  [^1] For power-law exponents 1/$21/2<\alpha\leq 3/$ 2, there is evidence [18, 20] that transport in models of this type crosses over to superdiffusion, which in turn gives way to unphysical instantaneous relaxation for $\alpha<1/2$. Indeed, the analysis in our paper is complicated for $\alpha\leq 3/2$ by the failure of Eq. (7) to converge. Restoring a finite-size cutoff in Eq. (8) and performing a numerical finite-size scaling analysis, not shown, we see behaviour consistent with the qualitative picture proposed in Ref. [18] and therefore with the absence of quasiballistic transport for $\alpha\leq 3/2$. to assume that $a(r),\,b(r)=o(r^{-3/2})$ as $r\to\infty$. We additionally assume that $a$ is non-zero (i.e. $a(r)\neq 0$ for some integer $r\neq 0$) to exclude Ising models without spin transport.
@@
By the continuity equation, the total current operator coincides with the rate of change of the “many-body position operator” [31], namely $\hat{J}=\sum_{n\in\mathbb{Z}}\hat{j}_{n}=\sum_{n\in\mathbb{Z}}n\partial_{t}\hat{S}^{z}_{n}$, where
@@
For generic, chaotic, inversion-symmetric Hamiltonians we expect diffusive spin transport above zero temperature (see Appendix A.2 for details). In such systems, a reasonable estimate for the infinite-temperature decay rate of the spin current can be obtained as follows [16, 29].
@@
where the function $D(u,v)=(u-v)a(u)a(v)+(u+v)a(u+v)(b(u)-b(v))$. At infinite temperature, this operator has zero mean but a non-zero, extensive variance, with a density
@@
In particular, minimizing $\tau_{\mathrm{eff}}^{-1}$ over the space of model parameters $a$ and $b$ should maximize the lifetime of the spin current and therefore lead to a maximal enhancement of quasiballistic spin transport [16]. We note that this strategy is justified provided $\hat{J}$ has negligible overlap with conserved quantities in the thermodynamic limit [32, 4, 33]. We will revisit this assumption more carefully below. First, we verify that our formula Eq. (8) is sufficient to capture the qualitative features of quasiballistic spin transport that have been observed numerically in long-range interacting XXZ chains [14, 15].
@@ III.1 Existence and uniqueness of an optimal anisotropy
Let us now specialize to spin chains with XXZ-type anisotropy, $b(r)=\Delta a(r)$. Then we can ask the following question: for a given choice of exchange interactions $a(r)$, is there a value $\Delta=\Delta^{*}$ of the model anisotropy that maximizes the lifetime of the spin current, as measured by $\tau_{\mathrm{eff}}$, and is this maximum unique? Existence and uniqueness of such a maximum is perhaps the simplest qualitative feature of the previous numerical studies [14, 15], and our first contribution is to show that this follows from our Eq. (8) on general grounds.
@@
Figure 1: Main figure: predicted optimal anisotropy $\Delta^{*}(\alpha)$ for quasiballistic spin transport as a function of the power-law exponent $\alpha$ (blue line). This is compared to the previous estimate $\Delta_{\mathrm{num}}^{*}(\alpha)\approx e^{2-\alpha}$ from numerics [14, 15] (red line). Both these loci intersect the integrable Haldane-Shastry point when $\alpha=2$ and recover the integrable XX point as $\alpha\to\infty$. Inset: tails of $\Delta^{*}(\alpha)$ versus $e^{2-\alpha}$ for large $\alpha$. The distinct asymptotic approach to the integrable XX point between the two estimates is clear.
@@ III.2 Power-law interactions
We now specialize further to power-law exchange interactions $a(x)=1/|x|^{\alpha}$, since these are both experimentally motivated [22, 23, 24, 25] and the subject of the previous numerical studies [14, 15] of quasiballistic spin transport. We further assume that $\alpha>3/2$ to ensure consistency of our approach. Then Eq. (11) predicts an exact expression for the optimal $\Delta^{*}$ for quasiballistic behaviour as a function of $\alpha$. The resulting locus in the space of model parameters is a ratio of two infinite series, namely
@@
This curve is plotted in Fig. 1, where it is compared to the previous estimate [14, 15] $\Delta^{*}_{\mathrm{num}}(\alpha)\approx e^{2-\alpha}$ for optimizing quasiballistic transport based on numerics. A nontrivial check on our prediction Eq. (12) is that just like the numerical estimate $\Delta^{*}_{\mathrm{num}}(\alpha)$, it “knows about” integrability. Specifically, Eq. (12) recovers the integrable Haldane-Shastry model, $\Delta^{*}(2)=1$. Similarly, Eq. (12) recovers the integrable XX model in the limit of nearest-neighbour interactions, $\Delta^{*}(\alpha)\to 0$ as $\alpha\to\infty$. Thus our analytical prediction recovers the two known integrable points in the space of power-law interacting XXZ models.
@@ III.3 Optimal relaxation rate and perturbative regime
Figure 2: Main figure: minimum decay rate for the spin current $\tau_{\mathrm{eff}}^{-1}(\Delta^{*}(\alpha))$ predicted by our approach (blue line), versus the relaxation rate extrapolated from the previous numerical estimate $\tau_{\mathrm{eff}}^{-1}(\Delta_{\textrm{num}}^{*}(\alpha))$ of the optimal anisotropy (red line). The decay rate vanishes as expected at the Haldane-Shastry point $\alpha=2$ and approaching the XX point as $\alpha\to\infty$. Inset: tails of $\tau_{\mathrm{eff}}^{-1}(\Delta^{*}(\alpha))$ and $\tau_{\mathrm{eff}}^{-1}(\Delta_{\textrm{num}}^{*}(\alpha))$ for large $\alpha$. It is apparent that both estimates decay exponentially in $\alpha$ at the same rate, with our estimate yielding a constant factor ($\approx 1/3$) improvement.
Finally, it is instructive to consider the optimal relaxation rate predicted by our approach. This is given by combining Eqs. (9) and (11) to obtain $\tau_{\mathrm{eff}}^{-1}(\Delta^{*}(\alpha))$, which is plotted in Fig. 2. As expected, this vanishes at $\alpha=2$ and as $\alpha\to\infty$. We further find that the relaxation rate remains small over the entire interval $2<\alpha<\infty$, consistent with the previous observation that the power-law XXZ chain is never “far” from ballistic transport [14].
While this is perhaps surprising for intermediate values of $\alpha$, it is expected that as $\alpha\to\infty$, power-law exchange interactions will merely yield a small (integrability-breaking) next-nearest-neighbour “hopping” correction $a(2)=2^{-\alpha}$ to the integrable XX point [14]. Moreover, for anisotropies $\Delta$ that are exponentially small in $\alpha$, the nearest-neighbour Ising “interaction” term $b(1)=\Delta$ will also be a small (integrability-preserving) perturbation to the XX point. This raises the questions of which small perturbation bottlenecks our estimate of the relaxation rate of the spin current for $\alpha\gg 1$ – on physical grounds, this should be the integrability-breaking hopping term – and how far our estimate for the optimal lifetime $\tau_{\mathrm{eff}}^{-1}(\Delta^{*}(\alpha))$ improves upon the lifetime $\tau_{\mathrm{eff}}^{-1}(\Delta^{*}_{\mathrm{num}}(\alpha))$ extrapolated from the previous numerical observations [14, 15] in this perturbative regime. (We note that the emergence of quasiballistic spin transport from these specific perturbations to the XX model was studied previously using different techniques [8].)
To explore these questions quantitatively, suppose that $\Delta\sim k\cdot c^{-\alpha}$ as $\alpha\to\infty$ for some constant $k$ and base $c>0$. Subject to this assumption, we find that the dominant contribution to the decay rate $\tau_{\mathrm{eff}}^{-1}$ depends sensitively on $c$, with
@@
as $\alpha\to\infty$. For $c\neq 2$, these formulas imply that the asymptotic relaxation rate is bottlenecked by the contribution from integrability-breaking next-nearest-neighbour hopping, as expected. However, for the specific value $c=2$, a nontrivial cancellation between hopping and interaction terms in Eq. (9) predicts a potentially slower relaxation rate $\tau^{-1}_{\mathrm{eff}}\sim\sqrt{\frac{(k-3)^{2}+1}{2}}\cdot 2^{-\alpha}$, which attains its minimum when $k=3$. It follows that $\tau^{-1}_{\mathrm{eff}}(\Delta^{*}_{\mathrm{num}}(\alpha))\sim\sqrt{5}\cdot 2^{-\alpha}$ is dominated by the hopping contribution, while $\tau^{-1}_{\mathrm{eff}}(\Delta^{*}(\alpha))\sim(1/\sqrt{2})\cdot 2^{-\alpha}$ by Eq. (13), predicting a decay rate for the spin current that is more than three times slower than would arise from next-nearest-neighbour hopping alone.

### v3

@@ I Introduction
Lattice models in physics generically exhibit diffusive transport of their local conserved charges at high temperature. The most commonly encountered exceptions to this rule arise in one spatial dimension, as a consequence of integrability. At one extreme, there exist models that can be mapped to non-interacting particles, such as the spin-$1/2$ XX [takahashi2005thermodynamics] and Haldane-Shastry [spinongas] chains, which therefore exhibit ballistic transport of all their local conserved charges [Antal, Sirker_2011, BulchandaniHa]. A more complicated intermediate behavior arises in the spin-$1/2$ XXZ chain, which can support ballistic, diffusive, or superdiffusive transport of spin, depending on the model anisotropy [Bertini].
@@
Thus, unlike in continuous systems, faster-than-diffusive transport in lattice models is generally linked to the presence of integrability [Castella_1995, Sirker_2011]. As one perturbs away from a given integrable point, such faster-than-diffusive dynamics usually gives way to normal diffusion. However, the relaxation timescales controlling this crossover to normal diffusion can be extremely long, even when the system under consideration is chaotic in all other respects [Ferreira_2020, DeNardis21, Roy_2023, McCarthy, McRoberts, wang2025breakdownsuperdiffusionperturbedquantum, Mierzejewski_2023, Yang_2024, HBS].
@@
In particular, several recent works have observed numerical evidence for long-lived and transiently ballistic (or “quasiballistic”) dynamics of conserved charges in systems that appear to be far from integrability. Throughout this paper, we follow Ref. [Mierzejewski_2023] in referring to this phenomenon as “quasiballistic transport”, with the understanding that this terminology refers to transiently ballistic dynamics rather than the specific behavior of any conventional linear-response transport coefficient. For example, spin transport in power-law-interacting spin-$1/2$ XXZ chains [Mierzejewski_2023, Yang_2024] (see also [Ferreira_2020]) and charge transport in interacting fermionic lattice models in any spatial dimension [HBS] can exhibit this phenomenon. Both sets of work focus on infinite temperature, where such behavior is intuitively least likely to occur. In the fermionic case, this physics was explained theoretically by computing an effective lifetime for the charge current, which could be made arbitrarily small as a function of the interaction range. Meanwhile, no detailed theoretical explanation has been proposed for quasiballistic spin transport in general and providing such an explanation is the primary goal of this work.
@@
The specific findings to be explained are as follows. It has been argued [Mierzejewski_2023, Yang_2024] that for spin-$1/2$ XXZ chains with power-law exchange interactions $J(r)\propto 1/r^{\alpha}$ between spins, long-lived and transiently ballistic modes of spin can arise for all values of the power-law exponent $\alpha$, with maximal enhancement of these ballistic modes for model anisotropies $\Delta^{*}(\alpha)\approx e^{2-\alpha}$ interpolating smoothly between the Haldane-Shastry chain (at $\alpha=2$) and the XX chain (as $\alpha\to\infty$), both of which exhibit ballistic spin transport as discussed above. However, the evidence that this locus exists and the extrapolation of its shape have so far been purely numerical. Our main contribution in this paper is an analytical prediction for the locus $\Delta^{*}(\alpha)$ on which the lifetime of the spin current is maximized, as a ratio of two infinite series. A nontrivial consistency check on our prediction is that it interpolates smoothly between the two ballistic points. However, our expression has the asymptotic behavior $\Delta^{*}(\alpha)\sim 3\cdot 2^{-\alpha}$ as $\alpha\to\infty$, whose exponential dependence on $\alpha$ matches the earlier conjecture from numerics [Mierzejewski_2023, Yang_2024], but decays more slowly when $\alpha$ is large.
@@
We emphasize that Hamiltonians of this type are not merely a theoretical curiosity. Recent advances in quantum devices have renewed interest in the dynamics of long-range interacting spin chains [Bohnet_2016, Schuckert20, Zu_2021, Joshi_2022, franke2023quantum]. Importantly, these platforms inherently favour the study of real-time dynamics. Moreover, their relatively high effective temperatures, which can be undesirable when it comes to exploring traditional questions about ground-state physics, are ideal for studying the far-from-equilibrium transport properties of interest in this work. There now exist several concrete proposals [arrazola2016digital, Bermudez17, Birnkammer22] for how to realize precisely the power-law interacting XXZ Hamiltonians studied here using trapped ions and a recent experiment [Kranzl23] probing their dynamics. It should, however, be borne in mind that relatively small power-law exponents $0\leq\alpha\leq 3$ are most physically natural in these platforms [fossfeig2024progresstrappedionquantumsimulation], while the discussion in our paper applies to $\alpha>3/2$.
@@
In this work, we pursue a less ambitious yet analytically tractable version of the path sketched above. Rather than ruling out ballistic spin transport fully for $\alpha\neq 2$, we rule out the possibility of ballistic spin transport protected by a strictly two-body conservation law, as it is at the Haldane-Shastry [haldane1994physicsidealsemiongas, Bernevig_2001, Sirker_2011] and XX points. Similarly, rather than attempting to evaluate the spin diffusion constant from the Kubo formula, we estimate the lifetime of the spin current directly from its instantaneous rate of change, a strategy that has proved effective in fermionic systems [HBS] (see also [Kim_2015]) and is reasonable in the absence of ballistic transport. We conclude by discussing the scope of our approach and the connection between integrability and transport more broadly. In particular, we prove that the only spin-$1/2$ chains of the form Eq. (1) with an exactly conserved spin current are Haldane-Shastry and XX chains (Theorem 1 of Appendix A.1).
@@ II Decay rate of the spin current
where $a(r)$ and $b(r)$ are arbitrary even functions and it will be convenient  [^1] For power-law exponents $1/2<\alpha\leq 3/2$, there is evidence [Schuckert20, Joshi_2022] that transport in models of this type crosses over to superdiffusion, which in turn gives way to unphysical instantaneous relaxation for $\alpha<1/2$. Indeed, the analysis in our paper is complicated for $\alpha\leq 3/2$ by the failure of Eq. (7) to converge. Restoring a finite-size cutoff in Eq. (8) and performing a numerical finite-size scaling analysis, not shown, we see behavior consistent with the qualitative picture proposed in Ref. [Schuckert20] and therefore with the absence of quasiballistic transport for $\alpha\leq 3/2$. to assume that $a(r),\,b(r)=o(r^{-3/2})$ as $r\to\infty$. We additionally assume that $a$ is non-zero (i.e. $a(r)\neq 0$ for some integer $r\neq 0$) to exclude Ising models without spin transport.
@@
By the continuity equation, the total current operator coincides with the rate of change of the “many-body position operator” [lieb1961two], namely $\hat{J}=\sum_{n\in\mathbb{Z}}\hat{j}_{n}=\sum_{n\in\mathbb{Z}}n\partial_{t}\hat{S}^{z}_{n}$, where
@@
For generic, chaotic, inversion-symmetric Hamiltonians we expect diffusive spin transport above zero temperature (see Appendix A.2 for details). In such systems, a reasonable estimate for the infinite-temperature decay rate of the spin current can be obtained as follows [HBS, Kim_2015].
@@
where the function D(u,v)=(uv)a(u)a(v)+(u+v)a(u+v)(b(u)b(v))D(u,v)=(u-v)a(u)a(v)+(u+v)a(u+v)(b(u)-b(v)). At infinite temperature, this operator has zero mean but a non-zero, extensive variance, with a density
@@
In particular, minimizing $\tau_{\mathrm{eff}}^{-1}$ over the space of model parameters $a$ and $b$ should maximize the lifetime of the spin current and therefore lead to a maximal enhancement of quasiballistic spin transport [HBS]. We note that this strategy is reasonable provided $\hat{J}$ has negligible overlap with conserved quantities in the thermodynamic limit [Zotos_1997, Sirker_2011, ProsenBound]. We will revisit this assumption more carefully below. First, we verify that our formula Eq. (8) is sufficient to capture the qualitative features of quasiballistic spin transport that have been observed numerically in long-range interacting XXZ chains [Mierzejewski_2023, Yang_2024]. Our approach is similar in spirit to the “moment method” for approximating the dynamical spin structure factor [de1958inelastic, redfield1968moment].
@@ III.1 Existence and uniqueness of an optimal anisotropy
Let us now specialize to spin chains with XXZ-type anisotropy, $b(r)=\Delta a(r)$. Then we can ask the following question: for a given choice of exchange interactions $a(r)$, is there a value $\Delta=\Delta^{*}$ of the model anisotropy that maximizes the lifetime of the spin current, as measured by $\tau_{\mathrm{eff}}$, and is this maximum unique? Existence and uniqueness of such a maximum is perhaps the simplest qualitative feature of the previous numerical studies [Mierzejewski_2023, Yang_2024], and our first contribution is to show that this follows from our Eq. (8) on general grounds.
@@
Figure 1: Main figure: predicted optimal anisotropy $\Delta^{*}(\alpha)$ for quasiballistic spin transport as a function of the power-law exponent $\alpha$ (blue line). This is compared to the previous estimate $\Delta_{\mathrm{num}}^{*}(\alpha)\approx e^{2-\alpha}$ from numerics [Mierzejewski_2023, Yang_2024] (red line). Both these loci intersect the integrable Haldane-Shastry point when $\alpha=2$ and recover the integrable XX point as $\alpha\to\infty$. Inset: tails of $\Delta^{*}(\alpha)$ versus $e^{2-\alpha}$ for large $\alpha$. The distinct asymptotic approach to the integrable XX point between the two estimates is clear.
@@ III.2 Power-law interactions
We now specialize further to power-law exchange interactions $a(x)=1/|x|^{\alpha}$, since these are both experimentally motivated [arrazola2016digital, Bermudez17, Birnkammer22, Kranzl23] and the subject of the previous numerical studies [Mierzejewski_2023, Yang_2024] of quasiballistic spin transport. We further assume that $\alpha>3/2$ to ensure consistency of our approach. Then Eq. (11) predicts an exact expression for the optimal $\Delta^{*}$ for quasiballistic behavior as a function of $\alpha$. The resulting locus in the space of model parameters is a ratio of two infinite series, namely
@@
This curve is plotted in Fig. 1, where it is compared to the previous estimate [Mierzejewski_2023, Yang_2024] $\Delta^{*}_{\mathrm{num}}(\alpha)\approx e^{2-\alpha}$ for optimizing quasiballistic transport based on numerics. A nontrivial check on our prediction Eq. (12) is that just like the numerical estimate $\Delta^{*}_{\mathrm{num}}(\alpha)$, it “knows about” integrability. Specifically, Eq. (12) recovers the integrable Haldane-Shastry model, $\Delta^{*}(2)=1$. Similarly, Eq. (12) recovers the integrable XX model in the limit of nearest-neighbour interactions, $\Delta^{*}(\alpha)\to 0$ as $\alpha\to\infty$. Thus our analytical prediction recovers the two known integrable points in the space of power-law interacting XXZ models. (Note that strictly speaking, $\alpha=\infty$ is an integrable line.)
@@ III.3 Optimal relaxation rate and perturbative regime
Figure 2: Main figure: minimum decay rate for the spin current $\tau_{\mathrm{eff}}^{-1}(\Delta^{*}(\alpha))$ predicted by our approach (blue line), versus the decay rate $\tau_{\mathrm{eff}}^{-1}(\Delta_{\textrm{num}}^{*}(\alpha))$ that we estimate from the previous numerical extrapolation of the optimal anisotropy (red line). The decay rate vanishes as expected at the Haldane-Shastry point $\alpha=2$ and approaching the XX point as $\alpha\to\infty$. Inset: tails of $\tau_{\mathrm{eff}}^{-1}(\Delta^{*}(\alpha))$ and $\tau_{\mathrm{eff}}^{-1}(\Delta_{\textrm{num}}^{*}(\alpha))$ for large $\alpha$. It is apparent that both estimates decay exponentially in $\alpha$ at the same rate, with our value of $\Delta^{*}(\alpha)$ yielding a constant-factor ($\approx 1/3$) reduction in the attainable decay rate as $\alpha\to\infty$.
Finally, it is instructive to consider the optimal relaxation rate predicted by our approach. This is given by combining Eqs. (9) and (11) to obtain $\tau_{\mathrm{eff}}^{-1}(\Delta^{*}(\alpha))$, which is plotted in Fig. 2. As expected, this vanishes at $\alpha=2$ and as $\alpha\to\infty$. We further find that the relaxation rate remains small over the entire interval $2<\alpha<\infty$, consistent with the previous observation that the power-law XXZ chain is never “far” from ballistic transport [Mierzejewski_2023].
While this is perhaps surprising for intermediate values of $\alpha$, it is expected that as $\alpha\to\infty$, power-law exchange interactions will merely yield a small (integrability-breaking) next-nearest-neighbour “hopping” correction $a(2)=2^{-\alpha}$ to the integrable XX point [Mierzejewski_2023]. Moreover, for anisotropies $\Delta$ that are exponentially small in $\alpha$, the nearest-neighbour Ising “interaction” term $b(1)=\Delta$ will also be a small (integrability-preserving) perturbation to the XX point. This raises the questions of which small perturbation bottlenecks (i.e. lower bounds) our estimate of the relaxation rate of the spin current for $\alpha\gg 1$ and very small values of $\Delta$ – on physical grounds, this should be the integrability-breaking hopping term – and how far our estimate for the optimal lifetime $\tau_{\mathrm{eff}}(\Delta^{*}(\alpha))$ improves upon the lifetime $\tau_{\mathrm{eff}}(\Delta^{*}_{\mathrm{num}}(\alpha))$ that we estimate from the previous numerical extrapolation [Mierzejewski_2023, Yang_2024] in this perturbative regime. (We note that the emergence of quasiballistic spin transport from similar perturbations to the XX model was studied previously using different techniques [Ferreira_2020].)
To explore these questions quantitatively, suppose that $\Delta\sim k\cdot c^{-\alpha}$ as $\alpha\to\infty$ for some constant $k$ and base $c>1$. Subject to this assumption, we find that the dominant contribution to the decay rate $\tau_{\mathrm{eff}}^{-1}$ depends sensitively on $c$, with
@@
as $\alpha\to\infty$. For suboptimal choices of $\Delta$ with $c\neq 2$, these formulas imply that $\tau_{\mathrm{eff}}^{-1}$ is lower-bounded by the contribution $\sim\sqrt{5}\cdot 2^{-\alpha}$ from integrability-breaking next-nearest-neighbour hopping, as expected. However, for the specific value $c=2$, a nontrivial cancellation between hopping and interaction terms in Eq. (9) predicts a potentially slower relaxation rate $\tau^{-1}_{\mathrm{eff}}\sim\sqrt{\frac{(k-3)^{2}+1}{2}}\cdot 2^{-\alpha}$, which attains its minimum when $k=3$. It follows that $\tau^{-1}_{\mathrm{eff}}(\Delta^{*}_{\mathrm{num}}(\alpha))\sim\sqrt{5}\cdot 2^{-\alpha}$ is dominated by the hopping contribution, while $\tau^{-1}_{\mathrm{eff}}(\Delta^{*}(\alpha))\sim(1/\sqrt{2})\cdot 2^{-\alpha}$ by Eq. (13), predicting a decay rate for the spin current that is more than three times slower than would arise from next-nearest-neighbour hopping alone, see Fig. 2.